OpenStudy (yanasidlinskiy):

Find all solutions of the equation: 2 sin (2A) +1

ganeshie8 (ganeshie8):

looks the equation is missing o.o

OpenStudy (yanasidlinskiy):

Whoops..There's "= 0" @ the end..

ganeshie8 (ganeshie8):

2sin(2A) + 1 = 0

ganeshie8 (ganeshie8):

subtract 1 both sides then divide 2

OpenStudy (yanasidlinskiy):

Umm..How would i do that?? I don't really understand..

ganeshie8 (ganeshie8):

2sin(2A) + 1 = 0 - 1 -1

ganeshie8 (ganeshie8):

2sin(2A) = -1

ganeshie8 (ganeshie8):

2sin(2A) = -1 /2 /2

ganeshie8 (ganeshie8):

sin(2A) = -1/2

OpenStudy (yanasidlinskiy):

Ohh.....Ok..That makes sense:D Now I can use your example for my other similar problems..Thank you:D

ganeshie8 (ganeshie8):

wait wait, we're not done yet

ganeshie8 (ganeshie8):

this is trigonometry problem, right ?

OpenStudy (yanasidlinskiy):

Yes.

ganeshie8 (ganeshie8):

we need to solve the unknown angle A

ganeshie8 (ganeshie8):

\(\large \sin(2A) = -1/2\) \(\large 2A = \arcsin (-1/2)\)

OpenStudy (yanasidlinskiy):

So, what would I do?? Multiply??

ganeshie8 (ganeshie8):

the answer is bit more complicated than that

ganeshie8 (ganeshie8):

if you're curious for what the final answer might looks like : \(\large A = \dfrac{7\pi}{12} + n\pi, ~\dfrac{11\pi}{12} + n\pi\)

OpenStudy (yanasidlinskiy):

Wow. I would never even would have thought about that..

ganeshie8 (ganeshie8):

sorry for weird looking expressions... but thats the final answer... we need to work a bit more to digest them ^

OpenStudy (yanasidlinskiy):

Ok.